OpenStudy (anonymous):

write y=6|x+3|-8 as a piecewise function, define √x^2

OpenStudy (anonymous):

y=f(x)=-6(x+3)-8 for x=(-∞,-3) = 6(x+3)-8 for x=(-3,∞) =-8 for x=-3

OpenStudy (anonymous):

sqrt(x^x)=|x|

OpenStudy (anonymous):

is this the answer? how did you do this?!

OpenStudy (anonymous):

@ZakaullahUET, sqrt(x^x)=|x| is wrong. you meant sqrt(x^2)=|x|

OpenStudy (anonymous):

yes yes, that was writing mistake

OpenStudy (anonymous):

@breeeee, draw the graph make it piecewise and write equation for every piece

OpenStudy (anonymous):

You get the piecewise function